OpenStudy (summersnow8):

In a section of horizontal pipe with a diameter of 3.00 cm the pressure is 100 kPa and water is flowing with a speed of 1.50 m/s. The pipe narrows to 2.00 cm. What is the pressure in the narrower region? Answer: ________ kPa

OpenStudy (fifciol):

continuity equation: \[A_1v_1=A_2v_2\]\[\pi \frac{ d_1^2 }{ 4 }v1=\pi \frac{ d_2^2 }{ 4 }v_2\] \[v_2=\frac{ d_1^2 }{ d_2^2 }v_1\] and Bernoulli equation: \[p_1 +\frac{ 1 }{ 2 }\rho_{water}v_1^2=p_2+\frac{ 1 }{ 2 }\rho_{water}v_2^2\] solve for p2

OpenStudy (fifciol):

99,54 kPa

OpenStudy (summersnow8):

\[p _{2}= \frac{ p _{1}v _{1^{2}} }{ v _{2^{2}} }\]

OpenStudy (summersnow8):

I dont think i have the equation right, because i am pretty sure the answer is 95.4, but i cant figure out how to get that

OpenStudy (fifciol):

oh , I made the mistake I chose 100 as density not 1000 as it should be I redo it and find 95,429 kPa

OpenStudy (fifciol):

\[p_2=p_1-\frac{ 1 }{ 2}\rho_{water}v_1^2[1-\left( \frac{ d_1 }{ d_2 } \right)^4]\]

OpenStudy (fifciol):

sorry it sholud be p1 + the rest

OpenStudy (fifciol):

so p2= 100 000 +500*2,25[1-5,0635]

OpenStudy (fifciol):

and then convert to kPa by dividing solution by 1000

OpenStudy (summersnow8):

wait, i didnt get that, can you show me what you plugged into p2=p1−12ρwaterv21[1−(d1d2)4]

OpenStudy (summersnow8):

\[p _{2}= (1000)+ .5 (1000)(1.5)^{2}(1-(3/4)^{4})\]

OpenStudy (summersnow8):

mean (3/2)^4

OpenStudy (fifciol):

yes

OpenStudy (fifciol):

that's correct

OpenStudy (summersnow8):

how come i get like -3570.31

OpenStudy (fifciol):

oh 100 kPa =100 000 Pa instead of 1000 put 100 000

OpenStudy (summersnow8):

ok